# Pretraining in Actor-Critic Reinforcement Learning for Robot Motion Control

### v1

@@ Abstract
The pretraining-finetuning paradigm has facilitated numerous transformative advancements in artificial intelligence research in recent years. However, in the domain of reinforcement learning (RL) for robot motion control, individual skills are often learned from scratch despite the high likelihood that some generalizable knowledge is shared across all task-specific policies belonging to a single robot embodiment. This work aims to define a paradigm for pretraining neural network models that encapsulate such knowledge and can subsequently serve as a basis for warm-starting the RL process in classic actor-critic algorithms, such as Proximal Policy Optimization (PPO). We begin with a task-agnostic exploration-based data collection algorithm to gather diverse, dynamic transition data, which is then used to train a Proprioceptive Inverse Dynamics Model (PIDM) through supervised learning. The pretrained weights are loaded into both the actor and critic networks to warm-start the policy optimization of actual tasks. We systematically validated our proposed method on seven distinct robot motion control tasks, showing significant benefits to this initialization strategy. Our proposed approach on average improves sample efficiency by 40.140.1% and task performance by 7.57.5%, compared to random initialization. We further present key ablation studies and empirical analyses that shed light on the mechanisms behind the effectiveness of our method.
@@ 1 Introduction
The pretraining-finetuning paradigm has enabled recent major breakthroughs in computer vision (He et al., 2022; Lu et al., 2019) and natural language processing (Devlin et al., 2019), most notably in the case of large language models (Touvron et al., 2023; Achiam et al., 2023). In the domain of robotics, a similar methodology with pre-initialization and fine-tuning has been explored in several works that integrate visual-language model (VLM) backbones for manipulation tasks (Brohan et al., ; Black et al., 2024; Team et al., 2024; 2025; Barreiros et al., 2025). However, these works only address the pretraining of the vision or language backbones, which have well-studied benefits and strategies, but do not directly address the robot embodiment information. While these imitation learning-based approaches offer good generalization to different tasks, they suffer from low-frequency execution and are primarily demonstrated on stable platforms and environments, rather than on dynamically unstable robotic platforms or under substantial external disturbances.
@@
In robot motion control, reinforcement learning (RL) with Proximal Policy Optimization (PPO) Schulman et al. (2017) has been used to successfully achieve a wide range of robust and agile motions (Hwangbo et al., 2019; Miki et al., 2022; Hoeller et al., 2023; Rudin et al., 2025; Choi et al., 2023; Zhang et al., 2025; Siekmann et al., 2021; Yang et al., 2023). However, skill acquisition is slow and resource-intensive because RL is generally sample-inefficient and each new task is typically learned tabula rasa, even within the same embodiment. Looking back at model-based control paradigms (Ferrolho et al., 2023; Sleiman et al., 2021; Bellicoso et al., 2019; Murphy et al., 2012), for a specific robot embodiment, there is knowledge that is sharable across solutions to different tasks, e.g., the joint kinematics and dynamics of the model. Our hypothesis is that warm-starting RL training in actor-critic architectures by incorporating such embodiment-aware knowledge into the initial model weights has the potential to improve policy performance and sample efficiency.
Several studies have proposed the development of a skill repertoire for robots (Hoeller et al., 2023) or the pretraining of low-level controllers with fine-grained skills (Peng et al., 2022; 2021), which a high-level controller can subsequently use. Our perspective differs significantly from these research efforts, as we do not require a dataset comprising expert-level skills or the retraining of the entire pipeline when adding a new skill. Furthermore, in these works, the final performance heavily relies on the quality of the learned skills and their relevance to the task at the fine-tuning stage, as they can not deviate significantly from behaviors in the original dataset.
@@
In contrast, this work presents a method for smart network initialization in the context of learning robot motion control with PPO, which outperforms the commonly used random initialization (He et al., 2015) across various tasks with the same embodiment. Our perspective on the problem is more fundamental, as we propose a task-agnostic approach that focuses solely on encapsulating embodiment-specific knowledge across tasks. We employ an exploration-based data collection strategy to systematically investigate states most likely to appear in the initial stages of the RL process, where the robot learns fundamental concepts about its embodiment, including limb kinematics, dynamics, and basic stability. With the collected data, we then train an embodiment-aware Proprioceptive Inverse Dynamics Model (PIDM). By initializing the actor-critic structure with the weights of the PIDM model, we provide the RL process with general knowledge from the initial stumbling stages of the vanilla training process, thus facilitating training. Consequently, our pretrained weights do not contain task-specific biases, but let them emerge naturally during RL training, as the entire network is updated in an end-to-end fashion. Our work serves as a user-friendly plug-in that does not require modifications to the established paradigm of motion control learning. We validate this approach on a diverse skill set for a quadrupedal robot, including pedipulation (Arm et al., 2024), locomotion (Mittal et al., 2023), and five parkour skills (Hoeller et al., 2023). Our approach improves performance by $\mathbf{7.5}$% and sample efficiency by $\mathbf{40.1}$%. In summary, our contributions are:
@@
  • A paradigm for weight initialization in robot motion control for RL, for one embodiment and across tasks, that improves performance and sample efficiency in the training process.
@@
  • Our approach is task-agnostic, i.e., applicable to various Partially Observable Markov Decision Process (POMDP) formulations involving different commands, observations, rewards, curricula, and terrains.
@@
  • Empirical validation of our proposed approach on numerous quadrupedal robot tasks showcases significant improvements in performance and efficiency.
@@ 2 Related Works
Pretraining in Reinforcement Learning (RL): Although RL excels on well-specified tasks, its limited sample efficiency remains a key challenge (Jin et al., 2021), which can be improved with pretraining. Xie et al. (2022) systematically summarized the efforts made to introduce the pretraining paradigm into RL, covering perspectives such as exploration, skill discovery, data coverage maximization, and representation learning. Among these, the works most related to ours are those that aim to pretrain good representations through exploratory pretraining with various objectives (Sekar et al., 2020; Curi et al., 2020; Nikolov et al., 2018). For example, Schwarzer et al. (2021) employ a combination of latent dynamics modelling and unsupervised goal-conditioned RL to pretrain useful representations that can be later fine-tuned to task-specific rewards. Allen et al. (2021) developed an approach to learn Markovian abstract states by combining inverse model estimation and temporal contrastive learning. Zheng et al. (2025) build a probabilistic model to predict which states an agent will visit in the future using flow matching and devise the RL update algorithm accordingly. While all the previously mentioned methods, similarly to our approach, use exploration to pretrain a model, there are a few fundamental differences. Firstly, many of these works often necessitate modifications to the RL update rule, whereas we aim to develop a drop-in technique that can be easily integrated into existing RL pipelines. Second, we address a more complex set of robotic tasks, incorporating high nonlinearity, noisy observation and massive domain randomization, as well as complex reward structures and environments. Finally, instead of trying to acquire complex behaviors, we focus on bootstrapping the initial learning steps across any task. Although driven by a similar aim, we find that none of these works constitute directly comparable baselines to our approach.
@@
Learning dynamics models via deep learning: Long et al. (2025) surveyed works on learning dynamics models from physical simulators, highlighting different model architectures and utilization strategies. Lutter & Peters (2023) further categorize such models by their reliance on prior knowledge, their degree of interpretability, and whether they enforce physical properties such as energy conservation. To address the sim-to-sim or sim-to-real gap of a trained policy in simulation, Christiano et al. (2016) propose computing what the simulation expects the resulting next state(s) will be, and then relying on a deep-learned inverse dynamics model to deduce the optimal action. This closely relates to our design of splitting the RL policy into a learned actor part and a pretrained inverse dynamics model. Learning dynamics for legged locomotion is challenging due to high non-linearities, sophisticated contact dynamics, and severe sensor noise from impacts. Levy et al. (2024) propose a semi-structured dynamics model consisting of a known a priori Lagrangian equation and an ensemble of learned external torque and noise estimators. Our approach makes no such assumptions, and we remain completely model-free. Xu et al. (2025) trained a neural simulator that is stable and accurate over a thousand simulation steps, utilizing a lightweight GPT-2 (Radford et al., 2019) architecture. In contrast, our architecture is significantly more compact and is also exposed to noise, domain randomization, and partial observability of the environment.
@@ 3 Preliminaries
Motion control problems are typically represented as Partially Observable Markov Decision Processes (POMDPs), where a policy $\pi:\mathcal{O}\rightarrow\mathcal{A}$ directly maps observations $\mathcal{O}$ to actions $\mathcal{A}$ and aims to maximize the cumulative reward. The reward function $R(s_{t},a_{t},s_{t+1})$ encodes task objectives, where $s_{t},s_{t+1}\in\mathcal{S}$ are the current and next state, respectively, and $a_{t}\in\mathcal{A}$ is the action taken at timestep $t$. Specific to robot motion control tasks, the observation is often the conjunction of command $\mathcal{C}$, proprioception $\mathcal{X}$, exteroception $\mathcal{X}_{e}$, and last action(s) $\mathcal{A}$.
In RL, a large family of actor-critic algorithms (Konda & Tsitsiklis, 1999) has been widely applied in robotics, among which Proximal Policy Optimization (PPO) (Schulman et al., 2017) is particularly prominent. These constitute an important class of RL algorithms that integrate policy optimization with value function estimation. The actor updates the policy that selects actions, while the critic estimates the value function of the current policy, thereby reducing variance and improving the stability of learning.
Existing works on RL (Lee et al., 2020; Miki et al., 2022; Vollenweider et al., 2022; Arm et al., 2024; Stolle et al., 2024; Portela et al., 2024; Sleiman et al., 2024) often parametrize both the actor and critic networks with a simple Multi-Layer Perceptron (MLP) and initialize its weights randomly (He et al., 2015). Due to the large variety of possible observation configurations, task and command specifications, and the diverse number of layers and input dimensions, pretraining a single model for all downstream tasks becomes impractical. We will address this by providing a modular network architecture and a well-defined pretraining task.
@@ 4.1 Problem formulation
Drawing inspiration from model-based control (Ferrolho et al., 2023; Sleiman et al., 2021; Bellicoso et al., 2019; Murphy et al., 2012), in robotics, the system’s target state $s_{t+1}$ is either known or usually easier to derive from the task formulation than the action $a_{t}$ necessary to get there. This is because $a_{t}$ is always dependent on the robot’s dynamics, which in RL are learned indirectly from experience through the simulator.
@@
We empirically demonstrate this in Section 5.2 and propose splitting the vanilla MLP structure into multiple distinct blocks (see Figure 3). One of these blocks is our proposed Proprioceptive Inverse Dynamics Model (PIDM), which we define as a mapping $I(a_{t}\mid x_{t-K:t+1},a_{t-K:t-1})$, where $x_{t}\in o_{t}$ denotes the proprioception at timestep $t$, $a_{t}$ denotes the action taken at timestep $t$, and $K$ denotes the length of the history sequence.
@@ 4.2 Overview
Our overall goal is to pretrain a PIDM model using supervised learning, which can later be integrated into the actor and critic networks of PPO. First, we collect transition data ($x_{t},a_{t},x_{t+1}$) in a task-agnostic manner from the RL training process of an exploration policy (Sekar et al., 2020). Important to note is that we solely collect transitions from the early stages of the RL training, rather than from expert policy rollouts for a specific task(s). On the one hand, this design ensures that the method does not rely on prior knowledge of the downstream tasks, nor on access to a (near) expert policy.
@@ 4.3 Exploration-based data collection
We employ an exploration-based data collection strategy, heavily inspired by previous works (Sekar et al., 2020; Curi et al., 2020; Nikolov et al., 2018; Chua et al., 2018), outlined in Figure 3. We use it to obtain data samples that capture the jittery, exploratory behaviors commonly observed in the early stages of RL. In practice, an exploration policy is trained with PPO, where the transitions from the on-policy rollouts are accumulated into a buffer. A probabilistic ensemble of PIDM models is frequently retrained using a bootstrap approach, where data is sampled with replacement from the buffer. The training of the exploration policy is primarily guided by the disagreement in predictions in the ensemble, as a measure of epistemic uncertainty for the PIDM inference. This incentivizes the policy to explore states where the accuracy of the PIDM can be improved with more data. Using the prediction error from a single PIDM model may seem attractive at first. However, we find that the resulting policy is prone to exploring only large-magnitude actions and high-frequency jittering, corresponding to the aleatoric uncertainty of the model. Secondary rewards added include a minimal set of regularizing rewards to constrain unwanted behaviors (e.g., high action rates, torques, or joint velocities) that are common to any task, as well as a term that rewards foot-air-time to encourage interaction with the terrain. During data collection, we employ standard domain randomization techniques for RL training (Miki et al., 2022; Lee et al., 2020; Kaidanov et al., 2024), such as varying the robot link masses, the friction coefficients, and applying random forces.
@@
Figure 3: Proprioceptive Inverse Dynamics Model (PIDM) architecture and its integration into the actor and critic networks. During pretraining of the PIDM, the dashed red parts of the network are included. However, when integrating into the actor-critic structure, those are removed and replaced by the encoder and decoder in purple.

### v2

# Pretraining in Actor-Critic Reinforcement Learning for Robot Locomotion

@@ Abstract
The pretraining-finetuning paradigm has facilitated numerous transformative advancements in artificial intelligence research in recent years. However, in the domain of reinforcement learning (RL) for robot locomotion, individual skills are often learned from scratch despite the high likelihood that some generalizable knowledge is shared across all task-specific policies belonging to the same robot embodiment. This work aims to define a paradigm for pretraining neural network models that encapsulate such knowledge and can subsequently serve as a basis for warm-starting the RL process in classic actor-critic algorithms, such as Proximal Policy Optimization (PPO). We begin with a task-agnostic exploration-based data collection algorithm to gather diverse, dynamic transition data, which is then used to train a Proprioceptive Inverse Dynamics Model (PIDM) through supervised learning. The pretrained weights are then loaded into both the actor and critic networks to warm-start the policy optimization of actual tasks. We systematically validated our proposed method with 9 distinct robot locomotion RL environments comprising 3 different robot embodiments, showing significant benefits of this initialization strategy. Our proposed approach on average improves sample efficiency by 36.936.9% and task performance by 7.37.3% compared to random initialization. We further present key ablation studies and empirical analyses that shed light on the mechanisms behind the effectiveness of this method.
@@ 1 Introduction
The pretraining-finetuning paradigm has enabled recent major breakthroughs in computer vision (he2022masked; lu2019vilbert) and natural language processing (devlin-etal-2019-bert), most notably in the case of large language models (touvron2023llama; achiam2023gpt). In the domain of robotics, a similar methodology with pre-initialization and fine-tuning has been explored in several works that integrate visual-language model (VLM) backbones for manipulation tasks (brohan_rt-2_nodate; black2024pi_0; team2024octo; team2025gemini; barreiros2025careful). However, these works only address the pretraining of the vision or language backbones, which have well-studied benefits and strategies, but do not endow robots with information about embodiment. While these imitation learning-based approaches offer good generalization to different tasks, they suffer from low-frequency execution and are primarily demonstrated on stable platforms and environments, rather than on dynamically unstable robotic platforms or under substantial external disturbances.
@@
In robot locomotion control, reinforcement learning (RL) with Proximal Policy Optimization (PPO) schulman2017proximal has been used to successfully achieve a wide range of robust and agile motions (hwangbo_learning_2019; miki_learning_2022; hoeller_anymal_2023; rudin_parkour_2025; beea57b6878542999db71105a7a8d1d9; zhang2025distillationpponoveltwostagereinforcement; siekmann2021blindbipedalstairtraversal; yang2023neuralvolumetricmemoryvisual). However, skill acquisition is slow and resource-intensive because RL is generally sample-inefficient and each new task is typically learned tabula rasa, even within the same embodiment. Looking back at model-based control paradigms (ferrolho2023roloma; full_mpc_sleiman; full_mpc_belliscoso; Murphy2012HighDD), for a specific robot embodiment, there is knowledge that is sharable across solutions to different tasks, e.g., the joint kinematics and dynamics of the model. Motivated by this, we posit that warm-starting RL training in actor-critic architectures by incorporating such embodiment-aware knowledge into the initial model weights has the potential to improve policy performance and sample efficiency.
@@
Our proposed method consists of three stages: exploration-based data collection, pretraining, and reinforcement learning. We first employ an exploration-based data collection strategy to systematically investigate states most likely to appear in the initial stages of the RL process, where the robot learns fundamental concepts about its embodiment, including limb kinematics, dynamics, and basic stability. With the collected data, we then train an embodiment-aware Proprioceptive Inverse Dynamics Model (PIDM). Finally, by initializing the actor-critic structure with the weights of the PIDM model, we provide the RL process with general knowledge from the initial stumbling stages of the vanilla training process, thus facilitate training. Our pretrained weights do not contain task-specific biases, but let them emerge naturally during RL training, as the entire network is updated in an end-to-end fashion.
There are a large body of works on offline-to-online reinforcement learning (ball_efficient_2023; hansen_td-mpc2_2024; nakamoto_cal-ql_2023) that aim to bootstrap online RL performance by utilization of a reward-labeled offline dataset. However, our method differ in the way that we aim to provide task-agnostic weights initialization for all possible downstream tasks of that specific embodiment. The unknownness and possible variation of downstream MDPs determine that it is impossible to include task-specific reward signal in the pretraining dataset, thus making the methods that require the target MDP to be fully known and free to explore in advance infeasible. Another line of studies have proposed the development of a skill repertoire for robots (hoeller_anymal_2023) or the pretraining of low-level controllers with fine-grained skills (peng_ase_2022; peng_amp_2021) that can be used by high-level controllers. Our perspective distinguish itself by the feature that we do not require a dataset comprising expert-level skills or the retraining of the entire pipeline when adding a new skill. Furthermore, in these works, the final performance heavily relies on the quality of the learned skills and their relevance to the task at the fine-tuning stage, as they can not deviate significantly from behaviors in the original dataset.
In contrast to aforementioned research, this work presents a method for smart network initialization in the context of learning robot locomotion with PPO, which outperforms the commonly used random initialization (he2015delving) across various tasks with the same embodiment. Our perspective on the problem is novel as we propose a task-agnostic approach that focuses solely on encapsulating embodiment-specific knowledge across tasks. It does not need reward signal of the task-specific downstream MDPs to be present in the pretraining dataset, and serves as a user-friendly plug-in that does not require modifications to the established paradigm of locomotion learning. We validate this approach with a diverse locomotion skill set and multiple robot embodiments consisting of two quadrupeds and one humanoid. Our approach improves performance by $\mathbf{7.3}$% and sample efficiency by $\mathbf{36.9}$%. The main contributions are:
  1. 1.
@@
    A paradigm of embodiment-specific weight initialization for RL in robot locomotion learning, that improves performance and sample efficiency in the training process.
  2. 2.
@@
    The initialization obtained this way is task-agnostic, i.e., applicable to various downstream Partially Observable Markov Decision Process (POMDP) formulations involving different commands, observations, rewards, curricula and terrains, as long as the same robot embodiment is retained.
  3. 3.
@@
    Extensive empirical validation of our proposed approach with various embodiments and tasks showcases significant improvements in performance and sample efficiency.
@@ 2 Related Works
Pretraining representations in RL Although RL excels on well-specified tasks, its limited sample efficiency remains a key challenge (jin2021bellmaneluderdimensionnew), which can be improved with pretraining. xie_pretraining_2022 systematically summarized the efforts made to introduce the pretraining paradigm into RL, covering perspectives such as exploration, skill discovery, data coverage maximization, and representation learning. The works most related to this are those that pretrain representations using an unlabeled (reward-free) offline dataset. schwarzer_pretraining_2021 employ a combination of latent dynamics modelling and unsupervised goal-conditioned RL to pretrain useful representations that can be later fine-tuned to task-specific rewards. allen_learning_2021 developed an approach to learn Markovian abstract states by combining inverse model estimation and temporal contrastive learning. zheng_intention-conditioned_2025 build a probabilistic model to predict which states an agent will visit in the future using flow matching, but it also necessitates the use of its own RL update algorithm and thus can not be used with existing prevalent RL algorithms. While all the previously mentioned methods operate on unlabeled datasets as ours does, there are a few fundamental differences. Firstly, we address a more complex set of robotic tasks, incorporating high nonlinearity, noisy observation and massive domain randomization, as well as complex reward structures and environments, and provide a demo supporting its sim-to-real transfer capability. Second, one single pretrained model for an embodiment is shown to be successfully transferred to multiple downstream tasks with various formulations of commands and observation spaces. Finally, instead of trying to exhaustively covering transitions that are possible in the environment during pretraining, we focus on bootstrapping the initial learning steps of downstream RL utilizing similar data given the high-dimensionality of robot locomotion tasks. That said, although driven by a similar aim, we find that none of these works constitute directly comparable baselines to our approach.
Learning dynamics models via deep learning
@@
long_survey_2025 surveyed works on learning dynamics models from physical simulators, highlighting different model architectures and utilization strategies. doi:10.1177/02783649231169492 further categorize such models by their reliance on prior knowledge, their degree of interpretability, and whether they enforce physical properties such as energy conservation. To address the sim-to-sim or sim-to-real gap of a trained policy in simulation, christiano_transfer_2016 propose computing what the simulation expects the resulting next state(s) will be, and then relying on a deep-learned inverse dynamics model to deduce the optimal action. This closely relates to our design of splitting the RL policy into a learned actor part and a pretrained inverse dynamics model. Learning dynamics for legged locomotion is challenging due to high non-linearities, sophisticated contact dynamics, and severe sensor noise from impacts. levy_learning_2024 propose a semi-structured dynamics model consisting of a known a priori Lagrangian equation and an ensemble of learned external torque and noise estimators. Our approach makes no such assumptions, and we remain completely model-free. xu2025neural trained a neural simulator that is stable and accurate over a thousand simulation steps, utilizing a lightweight GPT-2 (radford2019language) architecture. In contrast, our architecture is significantly more compact and is also exposed to noise, domain randomization, and partial observability of the environment.
Cross-task locomotion learning
AMP (peng_amp_2021) trains a policy that utilizes behaviors contained in the motion dataset to achieve the task objective, by combining task-rewards with style-rewards specified by an adversarial discriminator. ASE (peng_ase_2022) pretrains a low-level policy to map latent variables to behaviors depicted in the dataset, and later a task-specific high level policy is trained to specify latent variable for directing the low-level policy to accomplish the task goal. yang_multi-expert_2020 proposed multi-expert learning architecture (MELA), where they first train a set of experts with distinct skills, and then introduce a gating network which synthesize a weighted combination of experts and is finetuned jointly with the experts, resulting in an adaptive policy. rudin_parkour_2025 distill multiple terrain-specific expert policies into a single foundation policy via the DAgger(ross2011reduction) algorithm, which is then finetuned on a broader terrain set, and can be further finetuned on unseen terrain of test. All of them assume access to representations of high-utility skills, either in the form of a motion dataset or a set of expert task-specific policies. Close relation is expected between the task at runtime and behaviors of the experts/dataset, e.g. task at runtime can be solved by a combination of skills in those prior knowledge representations, or task at runtime is some task with domain shift (e.g. locomotion on harder terrain). In our work, we do not require access to expert skills directly related to the runtime task, but are interested in the formulation of universal knowledge for locomotion which can facilitate the learning of a possibly wide range of downstream tasks.
@@ 3 Preliminaries
Motion control problems are typically represented as Partially Observable Markov Decision Processes (POMDPs), where a policy $\pi:\mathcal{O}\rightarrow\mathcal{A}$ directly maps observations $\mathcal{O}$ to actions $\mathcal{A}$ and aims to maximize the cumulative reward. The reward function $R(s_{t},a_{t},s_{t+1})$ encodes task objectives, where $s_{t},s_{t+1}\in\mathcal{S}$ are the current and next state, respectively, and $a_{t}\in\mathcal{A}$ is the action taken at timestep $t$. Specific to robot motion control tasks, the observation is often the conjunction of command $\mathcal{C}$, proprioception $\mathcal{X}$, exteroception $\mathcal{X}_{e}$, and last action(s) $\mathcal{A}$. Compositions of these spaces are detailed in Appendix 7.
In RL, a large family of actor-critic algorithms (actor_critic_algo) has been widely applied in robotics, among which Proximal Policy Optimization (PPO) (schulman2017proximal) is particularly prominent. These constitute an important class of RL algorithms that integrate policy optimization with value function estimation. The actor updates the policy that selects actions, while the critic estimates the value function of the current policy, thereby reducing variance and improving the stability of learning.
Existing works on RL (lee2020learning; miki_learning_2022; vollenweider2022advanced; arm_pedipulate_2024; stolle_perceptive_2024; portela2025whole; sleiman_guided_2024) often parametrize both the actor and critic networks with a simple Multi-Layer Perceptron (MLP) and initialize its weights randomly (he2015delving). Due to the large variety of possible observation configurations, task and command specifications, and the diverse number of layers and input dimensions, pretraining a single model for all downstream tasks becomes impractical. We will address this by providing a modular network architecture and a well-defined pretraining task.
@@ 4.1 Problem formulation
Drawing inspiration from model-based control (ferrolho2023roloma; full_mpc_sleiman; full_mpc_belliscoso; Murphy2012HighDD), in robotics, the system’s target state $s_{t+1}$ is either known or usually easier to derive from the task formulation than the action $a_{t}$ necessary to get there. This is because $a_{t}$ is always dependent on the robot’s dynamics, which in RL is learned indirectly from experience through the simulator.
@@
We empirically demonstrate evidence supporting this hypothesis in Section 5.2 and propose splitting the vanilla MLP structure into multiple distinct blocks (see Figure 3). One of these blocks is our proposed Proprioceptive Inverse Dynamics Model (PIDM), which we define as a mapping $I(a_{t}\mid x_{t-K:t+1},a_{t-K:t-1})$, where $x_{t}\in o_{t}$ denotes the proprioception at timestep $t$, $a_{t}$ denotes the action taken at timestep $t$, and $K$ denotes the length of the history sequence.
@@ 4.2 Overview
Our overall goal is to pretrain a PIDM model using supervised learning, which can later be integrated into the actor and critic networks of PPO. First, we collect proprioceptive transition data ($x_{t},a_{t},x_{t+1}$) in a task-agnostic manner from the RL training process of an exploration policy (sekar2020planning). Important to note is that we solely collect transitions from the early stages of the RL training, rather than from expert policy rollouts for a specific task(s). On the one hand, this design ensures that the method does not rely on prior knowledge of the downstream tasks, nor on access to a (near) expert policy.
@@ 4.3 Exploration-based data collection
We employ an exploration-based data collection strategy, heavily inspired by previous works (pmlr-v97-pathak19a; sekar2020planning; curi2020efficient; nikolov2018information; chua_deep_2018), outlined in Figure 3. We use it to obtain data samples that capture the jittery, exploratory behaviors commonly observed in the early stages of RL. In practice, an exploration policy is trained with PPO, where the transitions from the on-policy rollouts are accumulated into a buffer. A probabilistic ensemble of PIDM models is frequently retrained using a bootstrap approach, where data is sampled with replacement from the buffer. The training of the exploration policy is primarily guided by the disagreement in predictions in the ensemble, as a measure of epistemic uncertainty for the PIDM inference. This incentivizes the policy to explore states where the accuracy of the PIDM can be improved with more data. Using the prediction error from a single PIDM model as intrinsic reward may seem probable and easier to implement at first sight. However, we find that its resulting policy is prone to exploring only large-magnitude actions and high-frequency jittering, corresponding to the aleatoric uncertainty of the model. Secondary rewards added include a minimal set of regularizing rewards to constrain unwanted behaviors (e.g., high action rates, torques, or joint velocities) that are common to any task, as well as a term that rewards foot-air-time to encourage interaction with the terrain. During data collection, we employ standard domain randomization techniques for RL training (miki_learning_2022; lee2020learning; kaidanov2024role), such as varying the robot link masses, the friction coefficients, and applying random forces.
@@
Figure 3: Proprioceptive Inverse Dynamics Model (PIDM) architecture and its integration into the actor network. During pretraining of the PIDM, the dashed red parts of the network are included. However, when integrating into the actor-critic structure, those are removed and replaced by the encoder and decoder in purple.